# Evolutive sandpiles

### v1

@@ Abstract
Sandpile group or Abelian sandpile model was the first example of a self-organized critical system studied by Bak, Tang and Wiesenfeld. The dynamics of the sandpiles occurs when the grains topple over a graph. In this study, we allow the graph evolve over time and change the topology at each stage. This results in the occurrence of phenomenons impossible in the classical sandpile models. Like, configurations over evolutive graphs that are always are unstable. We also experiment with the stabilization of configurations with a large number of grains at the center. This allow us to obtain fractals. Finally, we obtain some power laws associated with some evolutive graphs.
@@ 1 Introduction
The dynamics of the Abelian sandpile model, which was firstly studied by Bak, Tang and Wiesenfeld in [4], is carried out on a connected graph $G=(V,E)$ with an special vertex $q\in V$, called sink. Let $\mathbb{N}$ denote the set of non-negative integers and let $\mathbb{Z}$ denote the integer numbers. In the sandpile model, a configuration on $(G,q)$ is a vector ${c}\in\mathbb{N}^{V}$, in which entry ${c}_{v}$ is associated with the number of grains of sand (depending on the context it can also be used chips or dollars instead) placed on vertex $v$. The sink vertex is used to collect the sand getting out the system. Let $\widetilde{V}$ denote the set of non-sink vertices. Two configurations $c$ and $d$ are equal if $c_{v}=d_{v}$ for each non-sink vertex $v\in\widetilde{V}$. A non-sink vertex $v$ is called stable if ${c}_{v}$ is less than its degree $d_{G}(v)$, and unstable, otherwise. Thus, a configuration is called stable if every non-sink vertex is stable. The toppling rule in the dynamics of the model consists in selecting an unstable non-sink vertex $u$ and moving $d_{G}(u)$ grains of sand from $u$ to its neighbors, in which each neighbor $v$ receives $m_{(u,v)}$ grains of sand, where $m_{(u,v)}$ denote the number of edges between $u$ and $v$. Note toppling vertex $v_{i}$ in configuration ${c}$ corresponds to the subtraction the $i$-th row of the Laplacian matrix to ${c}$. Recall the Laplacian matrix $L(G)$ of a graph $G$ is such that the $(u,v)$-entry of $L(G)$ is defined as
@@
Starting with any unstable configuration and toppling unstable vertices repeatedly, we will always obtain [9, Theorem 2.2.2] a stable and unique configuration after a finite sequence of topplings. The stable configuration obtained from the configuration ${c}$ will be denoted by $s({c})$.
@@ 2 Evolutive sandpiles
Several different models of sandpiles have appeared, an account might be found at [6]. An novel extension of the sandpile models occur by changing the topology of the network over time, i.e., the edges of the graph are allowed to change over discrete-time, therefore the Laplacian matrix also will change over discrete-time. To make this more precise, let us introduce the evolutive Laplacian matrix $L(t)$ as follows. For $t\in\mathbb{N}$, the $u,v$-entry of the evolutive Laplacian matrix is given by
@@
where $f_{u,v}(t)$ is the function that provides the number of edges between vertices $u$ and $v$ at time $t$. Therefore, for this evolutive Laplacian matrix, there is associated an evolutive graph $G_{t}$ at time $t$ which has edge set $E_{t}$.
@@
Figure 1 shows the evolution of the graph over the time.
Then, the dynamics of this evolutive sandpile model starts at time $t=0$, with a configuration vector $c$ that provides the number of grain of sand at each vertex of $G_{0}$. A vertex $v\in G_{t}$ is unstable at time $t$ if $c_{v}\geq d_{G_{t}}(v)$. Then, at time $t$ each unstable non-sink vertex is toppled once, and the obtained configuration will be the configuration at the following time $t+1$. Note that the configuration obtained after toppling each unstable vertex of $G_{t}$ is unique, this is due to the commutativity of the toppling process see [9, Theorem 2.2.2].
Continuing with the previous example. Assume vertex $v_{3}$ is the sink vertex. Consider the initial configuration $c=(3,2)$. In Figure 2, we can observe an avalanche, that is, a sequence of topplings of unstable non-sink vertices and evolutions of the graph.
@@
As previously mentioned, in the clasic sandpile model, it is possible to show that for any finite graph $G$ with sink $s$, every configuration on $(G,s)$ eventually reaches a stable configuration under the toppling operations. However, over an evolutive graph with sink this is not always true. For instance, consider an evolutive graph, whose vertex set is $\{u,v,s\}$ where $s$ is the sink, that evolves on two stages where $E_{t}=E(G_{t})={uv,us}$ when $t$ is odd, and $E_{t}=E(G_{t})={uv,vs}$ when $t$ is even. Start at time $t=0$ with configuration $(1,0)$, where the first entry is associated with vertex $u$ and the second entry is associated with vertex $v$. Since $\deg_{G_{0}}(u)=1$, then $u$ is unstable and sends its only grain of sand to $v$. Then, at time $t=1$, the configuration is $(0,1)$, and vertex $v$ is unstable, since at this stage $\deg_{G_{1}}(v)=1$. Therefore, the configuration is $(1,0)$ when $t$ is even, and it is $(0,1)$ when $t$ is odd. Figure 3 shows the evolution of these configurations. Is is possible to obtain a criterion to decide whether a configuration over an evolutive graph would stabilize after a finite number of topplings?
@@ 3 Pattern formation
Many geometric structures have appeared within the context of sandpiles, most of them exhibiting interesting fractal patterns. One of the such geometric structures is obtained on the infinite square grid graph, where the starting configuration has a large number of grains at the center vertex $(0,0)$ and no grains elsewhere. After stabilizing this configuration, a fractal starts to emerge, see Fig. 4 (b). It turns out that the continuum scaling limit of this Abelian sandpile has the structure of an Apollonian circle packing, the details are out of the scope of this text for which the reader, interested in deeply understand the mathematics behind this fractal, is invited to see [10, 11]. Many variants have appeared since then, for example, several authors have studied the patterns obtained when the initial configuration has fixed height at each non-center vertex. In [12] it is investigated the geometry when the initial height is equal to 2 at the non-center vertices.
@@
We explore the behavior of the stabilization of a configuration over an evolutive graph in which a unique vertex is featured with a large number of grains of sand. In these experiments the evolutive graph has an infinite number of vertices, but the number of edges joining any pair of vertices is finite. First, we observe that classic fractals can be obtained by using the same graph over the time, see Figure 4 (a) and (b). In this figure, at (a) we see the stabilization of the all-zeros configuration except by 8 grains at the center of the infinite grid that remains the same over the time. And the stabilization occurs only on the 3-by-3 sub-grid for that reason we only show this sub-grid.. At (b), we stabilize the starting configuration with the all-zeros at each vertex except by 10,000,000 grains placed at the center of the infinite grid. At (c) and (d), the evolutive graph only have two stages, even cases correspond to the infinite square grid, and the odd cases correspond to the infinite grid in which each vertex at $(i,j)$ is adjacent only with vertices at sites $(i-1,j+1)$, $(i+1,j+1)$, $(i-1,j-1)$ and $(i+1,j-1)$. Note the graph in odd cases has two connected components. At (c), we see the stabilization of the all-zeros configuration except by 8 grains at the center over the described evolutive graph. And at (d), we have the stabilization of the configuration with the all-zeros at each vertex except by 500,000 grains placed at the center. It is interesting to observe that no fractal emerges in this evolutive sandpile, despite some symmetry over the horizontal and vertical axes might be identified. At (e) and (f), the evolutive graph only have two stages, even cases correspond to the infinite square grid adding additional vertical edges, and the odd cases correspond to the infinite square grid adding additional horizontal edges. In both cases, each note has degree 5. At (e), we see the stabilization of the all-zeros configuration except by 8 grains at the center over the described evolutive graph. And at (f), we have the stabilization of the configuration with the all-zeros at each vertex except by 5,000,000 grains placed at the center.
@@
Figure 4: (a) (c) and (e) are partial views of infinite evolutive graphs of stabilizations of a initial configuration having $8$ grains at the center evolving over time $t$. In (a) the evolutive graph is the same at any time. (b) The stabilization of the configuration with $10,000,000$ grains at the center over the same evolutive graph as in (a). In (c) and (e), the evolutive graphs evolve only over two different infinite graphs. (d) The stabilization of the configuration with $500,000$ grains at the center over same evolutive graph as in (c). (f) The stabilization of the configuration with $5,000,000$ grains at the center over same evolutive graph as in (e).
In Figure 5, we explore the evolutive graph consisting of two stages: the even case consists of the infinite graph in which vertex at site $(i,j)$ is adjacent only with vertices at sites $(i,j+1)$ and $(i,j-1)$, and the odd case consists of the infinite graph in which vertex at site $(i,j)$ is adjacent only with vertices at sites $(i+1,j)$ and $(i-1,j)$. The topological properties of this evolutive graph consists in that each vertex has degree two and there is an infinite number of connected components at each stage. At (a), we stabilize the configuration with only 8 grains at the center, this stabilization takes a total of 9 stages to get stabilized. At (b), we have the stabilization of the configuration with $400,000$ grains at the center. Changing the topology of the graph over time might result in new fractals and observations.
@@ 4 Power laws
Now, our focus lies on sandpile models which evolves over time, investigating whether it adheres to a power law similarly to stationary systems. Alan Kirman has proposed that economics could be effectively modeled as a network evolving over time, wherein agents have the ability to learn from past experiences with their neighbors, allowing links between them to either weaken or strengthen [8].
The process of the evolutive sandpile model starts by adding one grain at random to vertices, resulting in a configuration $c$ where each vertex contains at most one grain. The dynamics proceed as follows: At time $t$, one grain is added to a randomly selected vertex, leading to the stabilization of $G_{t}$. This stabilization results in a new configuration $c^{\prime}$, which serves as the input configuration for $G_{t+1}$, and this process repeats until the number of topplings becomes zero.
@@
The curves representing the number of topplings required to stabilize the evolutionary models are depicted in the left hand side of Figures 6 and 7 for $\mathcal{G}$ and $\mathcal{D}$, respectively. On the right-hand side, we depict the power law that best fits the probability distribution of our sample. In Figure 7(f), however, we considered not the probability distribution but rather the complementary cumulative probability distribution. The power laws were computed using the powerlaw package available on Python, see [2]. For model $\mathcal{G}$, the grids used for the simulations were of sizes $10$, $50$, and $100$. For model $\mathcal{D}$, the grid sizes were $50$, $100$, and $200$. While model $\mathcal{G}$ follows a power law, as the size of the grid increases in model $\mathcal{D}$, it becomes increasingly difficult to fit the model to a power law. This suggests that evolutive critical systems may not follow power laws.
@@
Figure 7: On the left-hand side, the number of topplings needed to stabilize model $\mathcal{D}$ is represented, based on 10,000 occurrences of different grid graph sizes: $10\times 10$, $50\times 50$, and $200\times 200$. On the right-hand side, the best power law fit to the distribution is displayed. The complementary cumulative probability was considered to plot Figure (7(f)).
@@ 5 Conclusions
Evolutive sandpile had the potential to enrich the classic sandpile models. It is worth to mention for instance that extension of Schelling segregation model over evolutive graphs in [7]. In this setting, it is possible to explore classic results and extend them to evolutive sandpiles. For example, within the algebraic perspective of the sandpiles. The sum of two configurations ${c}$ and ${d}$ is performed entry by entry. The sandpile sum ${c}\oplus{d}$ of two configurations is defined as $s({c}+{d})$. A configuration ${c}$ is recurrent if there exists a non-zero configuration ${d}$ such that ${c}={c}\oplus{d}$. Recurrent configurations play a central role in the dynamics of the Abelian sandpile model since recurrent configurations together with the sandpile sum form an Abelian group known as sandpile group [9, Chapter 4]. The sandpile group of $G$ is denoted by $K(G)$. One of the interesting features of the sandpile group of connected graphs is that the order $|K(G)|$ is equal to the number $\tau(G)$ of spanning trees of the graph $G$. For the reader interested in the mathematics of the sandpiles, we recommend the book [9] or the survey [1]. Is it possible to obtain such algebraic structure on non-trivial evolutive graphs. On the other hand, Baker and Norine introduced a dollar game in [5] which consists that the initial configuration assigns to each vertex $v$ an integer number of dollars. A vertex which has a negative number of dollars assigned to it is said to be in debt. Two moves are allowed, a vertex $v$ either borrow one dollar from each of its neighbors or give one dollar to each of its neighbors. The object of the game is to reach, through a sequence of moves, a winning configuration that is a configuration in which no vertex is in debt. A sequence of moves which achieves such a configuration is called a winning strategy. Baker and Norine developed a graph-theoretic analogue of the classical Riemann-Roch theorem which turn out in a characterization that allows to decide whether a winning strategy exists. Let $g=|E(G)|-|V(G)|+1$ and $N$ be the total number of dollars present at any stage of the game. Therefore, (1) if $N\geq g$, then there is always a winning strategy. (2) if $N<g-1$, then there is always an initial configuration for which no winning strategy exists. Is it possible to obtain winning strategies on this new setting?

### v2

@@ Abstract
The Abelian sandpile model was the first example of a self-organized critical system studied by Bak, Tang and Wiesenfeld. The dynamics of the sandpiles occur when the grains topple over a graph. In this study, we allow the graph to evolve over time and change the topology at each stage. This turns out in the occurrence of phenomena impossible in the classical sandpile models. For instance, configurations over evolutive graphs that are always unstable. We also experiment with the stabilization of configurations with a large number of grains at the center over evolutive graphs, this allows us to obtain interesting fractals. Finally, we obtain some power laws associated with some evolutive graphs.
Keywords: Evolutive sandpiles, fractals, power law
@@ 1 Introduction
The Abelian sandpile model was the first example of a self-organized critical system studied by Bak, Tang and Wiesenfeld. It was the first example of a self-organized critical system, which attempts to explain the occurrence of power laws in many natural phenomena [5] ranging on different fields like geophysics [20], optimization [9, 13, 2], economics [8] and neuroscience [10]. An exposure to self-organized-critically is provided in [5], and one for the sandpile model can be found in [18].
The dynamics of the Abelian sandpile model, which was first studied by Bak, Tang and Wiesenfeld in [6], is carried out on a connected graph $G=(V,E)$ with a special vertex $q\in V$, called sink. Let $\mathbb{N}$ denote the set of non-negative integers and let $\mathbb{Z}$ denote the integer numbers. In the sandpile model, a configuration on $(G,q)$ is a vector ${c}\in\mathbb{N}^{V}$, in which entry ${c}_{v}$ is associated with the number of grains of sand (depending on the context it can also be used chips or dollars instead) placed on vertex $v$. The sink vertex is used to collect the sand getting out of the system. Let $\widetilde{V}$ denote the set of non-sink vertices. Two configurations $c$ and $d$ are equal if $c_{v}=d_{v}$ for each non-sink vertex $v\in\widetilde{V}$. A non-sink vertex $v$ is called stable if ${c}_{v}$ is less than its degree $d_{G}(v)$, and unstable, otherwise. Thus, a configuration is called stable if every non-sink vertex is stable. The toppling rule in the dynamics of the model consists in selecting an unstable non-sink vertex $u$ and moving $d_{G}(u)$ grains of sand from $u$ to its neighbors, in which each neighbor $v$ receives $m_{(u,v)}$ grains of sand, where $m_{(u,v)}$ denote the number of edges between $u$ and $v$. Note toppling vertex $v_{i}$ in configuration ${c}$ corresponds to the subtraction the $i$-th row of the Laplacian matrix to ${c}$. Recall the Laplacian matrix $L(G)$ of a graph $G$ is such that the $(u,v)$-entry of $L(G)$ is defined as
@@
Starting with any unstable configuration and toppling unstable vertices repeatedly, we will always obtain [15, Theorem 2.2.2] a stable and unique configuration after a finite sequence of topplings. The stable configuration obtained from the configuration ${c}$ will be denoted by $s({c})$.
In this article we explore the behavior of the sandpile models on evolutive graphs, that is, graphs that change their topology over time. We call them evolutive sandpiles and define them in Section 2. This turns out in the occurrence of phenomena impossible in the classical sandpile models. For instance, configurations over evolutive graphs that are always unstable. In Section 3, we explore the patterns obtained from the stabilization of a starting configuration in which only one vertex contains grains. In the classical sandpile models, interesting fractals emerge from these experiments. And in Section 4, we investigate the power laws that better fit the number of topplings required to stabilize evolutive sandpile models.
@@ 2 Evolutive sandpiles
Several different models of sandpiles have appeared, an account might be found at [11]. A novel extension of the sandpile models occurs by changing the topology of the network over time, i.e., the edges of the graph are allowed to change over discrete-time, therefore the Laplacian matrix also will change over discrete time. To make this more precise, let us introduce the evolutive Laplacian matrix $L(t)$ as follows. For $t\in\mathbb{N}$, the $u,v$-entry of the evolutive Laplacian matrix is given by
@@
where $f_{u,v}(t)$ is the function that provides the number of edges between vertices $u$ and $v$ at time $t$. Therefore, for this evolutive Laplacian matrix, there is an associated an evolutive graph $G_{t}$ at time $t$ which has edge set $E_{t}$.
@@
Figure 1 shows the evolution of the graph over time.
Then, the dynamics of this evolutive sandpile model start at time $t=0$, with a configuration vector $c$ that provides the number of grain of sand at each vertex of $G_{0}$. A vertex $v\in G_{t}$ is unstable at time $t$ if $c_{v}\geq d_{G_{t}}(v)$. Then, at time $t$ each unstable non-sink vertex is toppled once, and the obtained configuration will be the configuration at the following time $t+1$. Note that the configuration obtained after toppling each unstable vertex of $G_{t}$ is unique, this is due to the commutativity of the toppling process see [15, Theorem 2.2.2].
Continuing the previous example, assume vertex $v_{3}$ is the sink vertex. Consider the initial configuration $c=(3,2)$. In Figure 2, we can observe an avalanche, that is, a sequence of topplings of unstable non-sink vertices and evolutions of the graph.
@@
As previously mentioned, in the classic sandpile model, it is possible to show that for any finite graph $G$ with sink $s$, every configuration on $(G,s)$ eventually reaches a stable configuration under the toppling operations. However, over an evolutive graph with a sink, this is not always true. For instance, consider an evolutive graph, whose vertex set is $\{u,v,s\}$ where $s$ is the sink, that evolves on two stages where $E_{t}=E(G_{t})={uv,us}$ when $t$ is odd, and $E_{t}=E(G_{t})={uv,vs}$ when $t$ is even. Start at time $t=0$ with configuration $(1,0)$, where the first entry is associated with vertex $u$ and the second entry is associated with vertex $v$. Since $\deg_{G_{0}}(u)=1$, then $u$ is unstable and sends its only grain of sand to $v$. Then, at time $t=1$, the configuration is $(0,1)$, and vertex $v$ is unstable, since at this stage $\deg_{G_{1}}(v)=1$. Therefore, the configuration is $(1,0)$ when $t$ is even, and it is $(0,1)$ when $t$ is odd. Figure 3 shows the evolution of these configurations. It is possible to obtain a criterion to decide whether a configuration over an evolutive graph would stabilize after a finite number of topplings?
@@ 3 Pattern formation
Many geometric structures have appeared within the context of sandpiles, most of them exhibiting interesting fractal patterns. One such geometric structure is obtained on the infinite square grid graph, where the starting configuration has a large number of grains at the center vertex $(0,0)$ and no grains elsewhere. After stabilizing this configuration, a fractal starts to emerge, see Fig. 4 (b). It turns out that the continuum scaling limit of this Abelian sandpile has the structure of an Apollonian circle packing, the details are out of the scope of this text for which the reader, interested in deeply understanding the mathematics behind this fractal, is invited to see [16, 17]. Many variants have appeared since then, for example, several authors have studied the patterns obtained when the initial configuration has a fixed height at each non-center vertex. In [19] it is investigated the geometry when the initial height is equal to 2 at the non-center vertices.
@@
We explore the behavior of the stabilization of a configuration over an evolutive graph in which a unique vertex is featured with a large number of grains of sand. In these experiments, the evolutive graph has an infinite number of vertices but the number of edges joining any pair of vertices is finite. First, we observe that classic fractals can be obtained by using the same graph over the time, see Figure 4 (a) and (b). In this figure, at (a) we see the stabilization of the all-zeros configuration except by 8 grains at the center of the infinite grid that remains the same over the time. Note that the stabilization occurs only on the 3-by-3 sub-grid for that reason we only show this sub-grid. At (b), we stabilize the starting configuration with the all-zeros at each vertex except by 10,000,000 grains placed at the center of the infinite grid. At (c) and (d), the evolutive graph only has two stages, even cases correspond to the infinite square grid, and the odd cases correspond to the infinite grid in which each vertex at $(i,j)$ is adjacent only with vertices at sites $(i-1,j+1)$, $(i+1,j+1)$, $(i-1,j-1)$ and $(i+1,j-1)$. Note the graph in odd cases has two connected components. At (c), we see the stabilization of the all-zeros configuration except by 8 grains at the center over the described evolutive graph. And at (d), we have the stabilization of the configuration with the all-zeros at each vertex except by 500,000 grains placed at the center. It is interesting to observe that no fractal emerges in this evolutive sandpile, however some symmetry over the horizontal and vertical axes might be identified. At (e) and (f), the evolutive graph only has two stages, even cases correspond to the infinite square grid adding additional vertical edges, and the odd cases correspond to the infinite square grid adding additional horizontal edges. In both cases, each note has degree 5. At (e), we see the stabilization of the all-zeros configuration except by 8 grains at the center over the described evolutive graph. And at (f), we have the stabilization of the configuration with the all-zeros at each vertex except by 5,000,000 grains placed at the center.
@@
Figure 4: (a) (c) and (e) are partial views of infinite evolutive graphs of stabilizations of an initial configuration having $8$ grains at the center evolving over time $t$. In (a) the evolutive graph is the same at any time. (b) The stabilization of the configuration with $10,000,000$ grains at the center over the same evolutive graph as in (a). In (c) and (e), the evolutive graphs evolve only over two different infinite graphs. (d) The stabilization of the configuration with $500,000$ grains at the center over the same evolutive graph as in (c). (f) The stabilization of the configuration with $5,000,000$ grains at the center over the same evolutive graph as in (e).
In Figure 5, we explore the evolutive graph consisting of two stages: the even case consists of the infinite graph in which vertex at site $(i,j)$ is adjacent only with vertices at sites $(i,j+1)$ and $(i,j-1)$, and the odd case consists of the infinite graph in which vertex at site $(i,j)$ is adjacent only with vertices at sites $(i+1,j)$ and $(i-1,j)$. The topological properties of this evolutive graph consist in that each vertex has degree two and there is an infinite number of connected components at each stage. At (a), we stabilize the configuration with only 8 grains at the center, this stabilization takes a total of 9 stages to get stabilized. At (b), we have the stabilization of the configuration with $400,000$ grains at the center. Changing the topology of the graph over time might result in new fractals and observations.
@@ 4 Power laws
Now, our focus lies on sandpile models in which adjacencies evolve over time, investigating whether they adhere to a power law similar to stationary systems. To accomplish this, we start the evolutive sandpile model by randomly assigning a non-negative number less than or equal to $\Delta(G_{0})-1$ to each vertex. This generates an initial random configuration. The dynamics proceed as follows: At time $t$, one grain is added to a randomly selected vertex, leading to the stabilization of $G_{t}$. This stabilization results in a new configuration $c^{\prime}$, which serves as the input configuration for $G_{t+1}$, and this process repeats until the number of topplings becomes zero.
@@
The curves representing the number of topplings required to stabilize the evolutionary models are depicted in the left-hand side of Figures 6 and 7 for $\mathcal{G}$ and $\mathcal{D}$, respectively. On the right-hand side, we depict the power law that best fits the probability distribution of our sample. In Figure 7(f), however, we considered not the probability distribution but rather the complementary cumulative probability distribution. The power laws were computed using the powerlaw package available on Python, see [3]. For model $\mathcal{G}$, the grids used for the simulations were of sizes $10$, $50$, and $100$. For model $\mathcal{D}$, the grid sizes were $50$, $100$, and $200$. A power law fitting is successful for both models, $\mathcal{G}$ and $\mathcal{D}$.
@@
Figure 7: On the left-hand side, the number of topplings needed to stabilize model $\mathcal{D}$ is represented, based on 10,000 occurrences of different grid graph sizes: $10\times 10$, $50\times 50$, and $200\times 200$. On the right-hand side, the best power law fit to the distribution is displayed.
@@ 5 Conclusions
Evolutive sandpile had the potential to enrich the classic sandpile models. It is worth mentioning for instance the extension of the Schelling segregation model over evolutive graphs in [12]. Also, Alan Kirman has proposed that economics could be effectively modeled as a network evolving over time, wherein agents have the ability to learn from past experiences with their neighbors, allowing links between them to either weaken or strengthen [14]. In this setting, it is possible to explore classic results and extend them to evolutive sandpiles.
@@
For example, within the algebraic perspective of the sandpiles, known as sandpile group. The sum of two configurations ${c}$ and ${d}$ is performed entry by entry. The sandpile sum ${c}\oplus{d}$ of two configurations is defined as $s({c}+{d})$. A configuration ${c}$ is recurrent if there exists a non-zero configuration ${d}$ such that ${c}={c}\oplus{d}$. Recurrent configurations play a central role in the dynamics of the Abelian sandpile model since recurrent configurations together with the sandpile sum form an Abelian group known as sandpile group [15, Chapter 4]. The sandpile group of $G$ is denoted by $K(G)$. One of the interesting features of the sandpile group of connected graphs is that the order $|K(G)|$ is equal to the number $\tau(G)$ of spanning trees of the graph $G$. For the reader interested in the mathematics of the sandpiles, we recommend the book [15] or the survey [1]. Is it possible to obtain such algebraic structure on non-trivial evolutive graphs.
On the other hand, Baker and Norine introduced a dollar game in [7] which consists of the initial configuration assigning to each vertex $v$ an integer number of dollars. A vertex that has a negative number of dollars assigned to it is said to be in debt. Two moves are allowed, a vertex $v$ either borrow one dollar from each of its neighbors or give one dollar to each of its neighbors. The object of the game is to reach, through a sequence of moves, a winning configuration that is a configuration in which no vertex is in debt. A sequence of moves that achieve such a configuration is called a winning strategy. Baker and Norine developed a graph-theoretic analog of the classical Riemann-Roch theorem which turn out in a characterization that allows one to decide whether a winning strategy exists. Let $g=|E(G)|-|V(G)|+1$ and $N$ be the total number of dollars present at any stage of the game. Therefore, (1) if $N\geq g$, then there is always a winning strategy. (2) if $N<g-1$, then there is always an initial configuration for which no winning strategy exists. Is it possible to obtain winning strategies in the evolutive setting?
Acknowledgement
The authors would like to thank Lothar Dirks for his helpful comments. The research of C.A. Alfaro is partially supported by the Sistema Nacional de Investigadores grant number 220797.
Appendix A Codes
In this appendix, we provide the codes that were used in the manuscript. The simulation of avalanche sizes, in addition to computing power-law distributions, was conducted using the Python programming language.
Listing 1: Evolutive function
import numpy
def evolutive(G: dict, c: list):
L = list()
t = 0
for i in range(10000):
avalanche_size = 0
random_vertex = numpy.random.randint(1, len(G[t].keys()))
c[random_vertex] += 1
while True:
c, k = stabilize(G[t],c)
avalanche_size += k
t += 1
t %= len(G.keys())
if k == 0: break
L.append(avalanche_size)
return L
The following demonstrates the use of the stabilize function to stabilize sandpiles on general graphs, along with the incorporation of the auxiliary topple method for toppling vertices.
Listing 2: Stabilization
def topple(G: dict, c: list, v: int):
n = len(G[v])
grains = c[v]
if c[v] >= n:
for w in G[v]:
c[w] += 1
c[v] -= 1
return grains >= n
def stabilize(G: dict, c: list):
k = 0
for v in range(1,len(G.keys())):
is_topple = topple(G,c,v)
if is_topple: k += 1
return c, k